A body is 2.94kg weight on moon . Then what is his actual weight ?

Grade:9

4 Answers

Aayush Raman
6 years ago
W = (Mg)/6
2.94=M*9.8/6=M*1.63
 
THEREFORE M= 2.94/1.63=1.8 APPROX
ACTUAL WEIGHT = MG = 1.8*9.8 = 17.6 APPROX
 
C.Vishaal
6 years ago
The weight on the moon is just 1/6 that of the weight on earth.
 
Therefore, it is 2.94 x 9.8 / 6
                                            = 4.802 approx.
Sanju
6 years ago
Acceleration due to gravity on the surface of the moon, g= 1.625 m/s
Given that m*g= 2.94 kg m/s
Therefore, mass, m of the body is: 2.94 kg m/s  / g= 2.94 kg m/s / 1.625 m/s= 1.809 kg
So the actual weight = mg, where g is the acceleration due to gravity on the surface of the earth.
ACTUAL WEIGHT = m*g = 1.809 kg * 9.81 m/s2 = 17.75 kg m/s
ayush
6 years ago
it is 2.94 kg on moon   GIVEN
the body on earth is 6times then at of moon 
 
2.94 * 6    WEIGHT 
 
FOR MASS   WE KNOW    F=W = mg  so g on earth is 9.8m/s^2      
 
 
so 2.94*6/9.8